OpenStudy (anonymous):

Use f(x) = 3x − 5 and g(x) = 4 − x2 to evaluate the expression. (a) (f o g)(−2) (g o f)(−2)

OpenStudy (kainui):

So basically f o g means just take the function f(x) and just take all the places where there's an x and put in whatever g(x) is. So: f(x)=3x-5 f(x)=3(g(x))-5 f(x)=3(4-x^2)-5 now you can plug in -2 to it like they ask you, after a little more simplification if you so desire. I can help you a little more if you have any questions.